OpenStudy (anonymous):

Solve. 12/n=48/100 n =

OpenStudy (anonymous):

@horsegirl27 @satellite73

OpenStudy (horsegirl27):

I will help. @satellite73 could you too?

OpenStudy (anonymous):

i dont think he is here

OpenStudy (horsegirl27):

12/25*12/n I think is how to set it up.

OpenStudy (horsegirl27):

You try to solve

OpenStudy (horsegirl27):

I mean 12/25=12/n

OpenStudy (anonymous):

n=25

OpenStudy (horsegirl27):

12/n=12/48 12n=576(divide both sides by 12) n=48